A man who weighs 200 lbs stands 40 ft from one end of a uniform 200-ft scaffold. A 60 lb pile of bricks has its center of gravity 70 ft from the same end, and the scaffold weighs 90 lbs. If the scaffold supported at the two ends, find the force on each support.
1
Expert's answer
2012-08-17T07:50:05-0400
Let F1 is the force on first support (at the end where a man stands)
and F2
is the force on second support.
200*160*g + 60*130*g + 90*100*g = F1 *
200
F1 = 224*g = 996.76 N
200*40*g + 60*70*g + 90*100*g = F2 * 200
F2 =
106*g = 471.68 N
Please note that g = 9.81 N/kg, 1 lb = 0.4536 kg.